Beckheart is seeking fi nancing for its inventory. Safe-Proof
Warehouses off ers space in their facility for Beckheart’s inventory.
They off er loans with a 15 percent APR equal to 60 percent of the
inventory. Monthly fees for the usage of the warehouse are $500 plus
0.5 percent of the inventory’s value. If Beckheart has saleable inventory
of $2 million, answer the following:
a. How much money can the fi rm borrow?
b. What is the interest cost of the loan in dollars over a year?
c. What is the total amount of fees to be paid in a year?
d. What is the eff ective annual rate of using Safe-Proof to fi nance
Beckheart’s inventory?
1
Expert's answer
2019-11-28T11:37:54-0500

Beck heart's inventory


We need to find the borrowed money, interest cost of the loan, total amount of fee and effective annual rate


Solution:


Beck heart has saleable inventory of $2 million


a).


money borrowed by firm = 60 % of $2 million

"=\frac {60} {100} \times 2000000 = \$1200000"

b).


Interest cost of the loan over a year = 15% of 1200000


"= \frac {15}{100} \times 1200000 = \$180000"

c).


Total amount of fees to be paid in a year =

"=(12 \times 500) + ( 2000000 \times 0.5 \times 12)"



"= 6000 + 120000 = \$ 126000"

d).


Effective annual rate of using safe - proof to finance Beck heart's inventory


"=\frac {\$180000 + 126000} {2000000} \times 100 \% = 15.3\%"